At equilibrium, the concentrations of reactants and products can be predicted using the equilibrium constant, KcKcK_c, which is a mathematical expression based on the chemical equation. For example, in the reaction

aA+bB⇌cC+dDaA+bB⇌cC+dD

where aaa, bbb, ccc, and ddd are the stoichiometric coefficients, the equilibrium constant is

Kc=[C]c[D]d[A]a[B]bKc=[C]c[D]d[A]a[B]b

where [A][A], [B][B], [C][C], and [D][D] are the equilibrium concentrations. If the reaction is not at equilibrium, the quantity can still be calculated, but it is called the reaction quotient, QcQcQ_c, instead of the equilibrium constant, KcKcK_c.

Qc=[C]tc[D]td[A]ta[B]tbQc=[C]tc[D]td[A]ta[B]tb

where each concentration is measured at some arbitrary time ttt.

Part A

A mixture initially contains AA, BB, and CC in the following concentrations: [A][A]A_1 = 0.450 MM , [B][B]B_1 = 0.750 MM , and [C][C]C_1 = 0.650 MM . The following reaction occurs and equilibrium is established:

A+2B⇌CA+2B⇌C

At equilibrium, [A][A]A_2 = 0.300 MM and [C][C]C_2 = 0.800 MM . Calculate the value of the equilibrium constant, KcKcK_c.

Express your answer numerically.
Answer #1

A + 2B ⇌    C

I(M) 0.450 0.750 0.650

E(M) 0.450-x 0.750-2x 0.650+x

According to question,

0.650+x = 0.800

x = 0.150 M

[A]eq = 0.450-x = 0.300 M

[B]eq = 0.750-2x= 0.450 M

[C]eq = 0.650+x = 0.800 M

Kc = [C]eq/[A]eq[B]2eq

Kc = [0.800 M/0.300 M x (0.450 M)2]

Kc = 13.17